OpenStudy (anonymous):

Find the difference quotient for f(x)=5-7x+8x^2

OpenStudy (isaiah.feynman):

Know the difference quotient formula?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

f(a+h)-f(a)/h

OpenStudy (isaiah.feynman):

Oh lets use this one f(x+h)-f(x)/h, since we don't have a.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I already found f(a) and f(a+h)

OpenStudy (isaiah.feynman):

Okay can I see them please.

OpenStudy (anonymous):

f(a)=5-7a+8a^2 f(a+h)=5-7(a+h)+8(a+h)^2

OpenStudy (isaiah.feynman):

No, doesn't look correct. "a" is used to denote a special number that you must find the derivative at. We don't have that special number so we can't just change x to a. That's why I said we should the formula I gave.

OpenStudy (anonymous):

It is correct because I checked the answer and it is right

OpenStudy (anonymous):

I will figure it out, thanks for the help

OpenStudy (isaiah.feynman):

Okay...if you can use x or a then we would get the same answer.

OpenStudy (isaiah.feynman):

So did you get 16a+8h-7? @hannahmarie620

OpenStudy (anonymous):

for what?

OpenStudy (isaiah.feynman):

For the difference quotient.

OpenStudy (anonymous):

no

OpenStudy (isaiah.feynman):

What did you get?

OpenStudy (anonymous):

wait sorry yes I did get that

OpenStudy (anonymous):

I forgot to cancel the h's out

OpenStudy (isaiah.feynman):

Good! I have a confession I got 16x+8h-7. I still used x instead of a. :P